### v3

@@ Abstract
In this paper, we address a manifold constrained nonsmooth optimization problem involving the composition of a weakly convex function and a smooth mapping. To find a stationary point of the target problem, we propose a variable smoothing-type algorithm by combining the ideas of (i) translating the constrained problem into a Euclidean optimization problem with a smooth parametrization of the constraint set; (ii) exploiting a sequence of smoothed surrogate functions, of the cost function, given with the Moreau envelope of a weakly convex function. The proposed algorithm produces a vector sequence by the gradient descent update of a smoothed surrogate function at each iteration. In a case where the proximity operator of the weakly convex function is available, the proposed algorithm does not require any iterative solver for subproblems therein. By leveraging tools in the variational analysis, we show the so-called gradient consistency property, which is a key ingredient for smoothing-type algorithms, of the smoothed surrogate function used in this paper. Based on the gradient consistency property, we also establish an asymptotic convergence analysis for the proposed algorithm together with convergence rate analysis. Numerical experiments demonstrate the efficacy of the proposed algorithm.
@@ 1 Introduction
Problem 1.1 has been serving as key models, e.g., for finding sparse solutions, in wide range of signal processing and machine learning. Such applications include, e.g., sparse Principal Component Analysis (PCA) [1], sparse variable PCA [2], robust sparse PCA [3], orthogonal dictionary learning/robust subspace recovering [4], and sparse spectral clustering [5].
@@
One of simple cases of Problem 1.1 is a convex case where $h$ and $g$ are convex, $\mathfrak{S}$ is linear, and $C\coloneqq\mathcal{X}$. For the convex case, the so-called proximal splitting algorithms have been extensively applied as computationally reliable algorithms (see, e.g., [6, 7, 8]). The algorithms consist of simple steps involving the gradient of $h$ and the so-called proximity operator $\mathrm{prox}_{\mu g}:\mathcal{Z}\to\mathcal{Z}:\widebar{\bm{z}}\mapsto=\mathop{\mathrm{argmin}}\limits_{\bm{z}\in\mathcal{Z}}\left(g(\bm{z})+\frac{1}{2\mu}\left\lVert\bm{z}-\widebar{\bm{z}}\right\rVert^{2}\right)$ of $g$ with index $\mu>0$. Although the computation of $\mathrm{prox}_{\mu g}$ requires, in general, to solve a convex optimization problem, the proximity operators for many commonly-used functions, e.g., $\ell_{1}$-norm and nuclear norm, have closed-form expressions (see, e.g., [6, 9]).
@@
In Problem 1.1, restricting choices of $g$ within convex functions are not always sufficient [10, 11, 12, 13, 14, 15] in order to enhance estimation performance especially for sparsity-aware applications. For example, $\ell_{1}$-norm has been widely used as a best convex approximation of the naive sparsity promoting function, i.e., $\ell_{0}$-pseudonorm, which counts the number of nonzero entries of a vector. However, the use of $\ell_{1}$-norm as $g$ leads to some underestimation of the desired sparse target because $\ell_{1}$-norm is a coercive function (see, e.g., [16]), i.e., $\ell_{1}$-norm excessively penalizes large magnitude of target. To overcome the underestimation effect caused mainly by the convex sparse promoting function, the utilization of weakly convex functions has been attracting great attention [10, 12]. Recent advances in proximal splitting techniques [11, 13, 14, 15] encourage us to find a global minimizer of Problem 1.1 under the convexity of $h+g\circ\mathfrak{S}$ with a linear operator and even with a weakly convex $g$.
Table 1: Summary of existing algorithms for Problem 1.1 and the proposed algorithm.
Method Ref. $g$ $\mathfrak{S}$ $C$ Single loop ?
Proximal linear [17] convex smooth unconstrained No
[18] weakly convex smooth unconstrained No
[19] convex smooth manifold No
Proximal gradient [20] convex linear manifold No
Subgradient [21] weakly convex identity manifold Yes
Variable smoothing [22] convex linear unconstrained Yes
[23] weakly convex linear unconstrained Yes
[24] convex linear manifold Yes
[25] weakly convex linear manifold Yes
Alg. 1 weakly convex smooth manifold Yes
@@
Beyond the convex cases, nonconvex optimization problems with nonconvex constraint have a great potential for flexible formulation to align with real-world tasks. Indeed, the minimization of nonconvex $g\circ\mathfrak{S}$ with a nonlinear smooth mapping $\mathfrak{S}$ is known to have extremely wide applications including, e.g., robust phase retrieval [26], nonnegative matrix factorization [27], robust matrix recovery [28], and robust blind deconvolution [29]. In particular, manifold constrained optimization has been continuingly attracting great attention mainly by its remarkable expressive ability of the target conditions expected to be achieved by a solution (see, e.g., [30, 31, 32, 33]). Hence, Problem 1.1 is certainly of our great interest.
Literature review
We review some related works for finding a stationary point, of Problem 1.1, which satisfies a certain necessary condition called a first-order optimality condition (see (25)) to be a local minimizer. Table 1 summarizes distinctions of target settings among algorithms in the related works.
Proximal-type algorithms. The proximal gradient method (also known as an instance of the forward-backward splitting method, e.g., [6, 8]) has been extended by [20] from the unconstrained setting to the setting of Problem 1.1 with a special case where $C$ is an embedded submanifold of $\mathcal{X}$, $g$ is convex and $\mathfrak{S}$ is linear. Such an extended proximal gradient method [20] requires some iterative solvers for a certain proximal subproblem at every iteration, which is defined as convex optimization problems over the tangent (vector) space (see (89)). So far, closed-form solutions of such proximal subproblems have not yet been found even for simple cases, e.g., $g\coloneqq\left\lVert\cdot\right\rVert_{1}$. For Problem 1.1 where $\mathfrak{S}$ is smooth but $g$ remains to be convex, the prox-linear method [18, 17] for the unconstrained setting has also been extended by [19] to the manifold constrained setting. Under a certain condition, e.g., the Lipschitz continuity of $\mathrm{D}\mathfrak{S}(\cdot)$ (see (33)), the prox-linear methods [18, 17, 19] are guaranteed to converge to a stationary point by passing through a subsequence of the generated sequence. Moreover, for the prox-linear methods, the iteration complexity to obtain an $\epsilon$-approximate stationary point is given by $\mathcal{O}(\epsilon^{-2})$, where $\mathcal{O}$ stands for the Landau’s Big O notation. However, these prox-linear methods require some iterative solver for solving a convex optimization problem to update estimates $(\bm{x}_{n})_{n=1}^{\infty}\subset\mathcal{X}$.
@@
In short, every update in the above proximal-type algorithms applicable for Problem 1.1 heavily relies on iterative solvers for solving subproblems. Convergence analyses of such proximal-type algorithms have been established basically with an exact solution to the subproblem, although such an exact solution is not available in general within a finite number of iterations of an iterative solver. Moreover, an iterative solver sometimes takes a lot of time to obtain even an approximate solution to a subproblem as will be shown in numerical experiments (see Section 5.2). These situations encourage us to develop a single-loop algorithm, such that any iterative solver is not required for solving subproblems at every iteration, for Problem 1.1.
Smoothing-type algorithms. For finding a stationary point of a nonsmooth function $J:\mathcal{X}\to\mathbb{R}$, smoothing methods have been extensively developed [34, 35, 36, 37, 38, 39, 40] with a smoothed surrogate function, say $J^{\langle\mu\rangle}:\mathcal{X}\to\mathbb{R}\ (\mu\in\mathbb{R}_{++})$, having the (sub-)gradient consistency property (see Remark 4.5) with $\lim_{\mathbb{R}_{++}\ni\mu\to 0}J^{\langle\mu\rangle}(\bm{x})=J(\bm{x})\ (\bm{x}\in\mathcal{X})$, where the gradient consistency property has been utilized as a key tool to express the first-order optimality condition in terms of a gradient sequence of $\nabla J^{\langle\mu\rangle}$ with $\mu\searrow 0$. The basic update of a smoothing gradient method [35] is given by $\bm{x}_{n+1}\coloneqq\bm{x}_{n}-\gamma_{n}\nabla J^{\langle\mu_{n}\rangle}(\bm{x}_{n})\ (n\in\mathbb{N})$ with a properly chosen stepsize $\gamma_{n}\in\mathbb{R}_{++}$, where smoothing parameters $\mu_{n}$ are conditionally updated with $\sigma\in(0,1)$ and $\delta\in\mathbb{R}_{++}$ by $\mu_{n+1}\coloneqq\sigma\mu_{n}$ if $\left\lVert\nabla J^{\langle\mu_{n}\rangle}(\bm{x}_{n})\right\rVert<\delta\mu_{n}$; and $\mu_{n+1}\coloneqq\mu_{n}$ otherwise. Thanks to the (sub-)gradient consistency property of $J^{\langle\mu\rangle}$, a cluster point of $(\bm{x}_{n})_{n=1}^{\infty}$ is guaranteed to be a (Clarke) stationary point of $J$ [35]. Based on the conditional update of smoothing parameters, many numerical algorithms, e.g., accelerated gradient methods [37, 40], projected gradient methods [34, 36] and manifold constrained smoothing methods [38, 39], for updating $(\bm{x}_{n})_{n=1}^{\infty}$ have been established with asymptotic convergence analyses in terms of a stationary point (or the optimal value of the cost function for convex optimization). In contrast, convergence rate analyses of smoothing methods have not been found except for a few papers, e.g., [41, 42] for an optimization problem involving $\ell_{p}$-quasinorm $(0<p\leq 1)$ (see Remark 4.10).
Another class of smoothing-type methods has been proposed with iteratively decreasing smoothing parameters $(\mu_{n})_{n=1}^{\infty}$ [22, 43, 44, 23, 45, 46, 24, 25, 47], where these methods are called variable/adaptive/dynamic smoothing methods. Thanks to the use of iteratively decreasing smoothing parameters, convergence rate analyses have been established for variable smoothing methods. For convex optimization problems, convergence rates in terms of the optimal value of the cost function have been shown for variable smoothing algorithms [22, 43, 44, 45, 46, 47], where an exceptional paper [46] guarantees that a generated sequence converges to a minimizer of the cost function.
Recently, for nonconvex nonsmooth optimization problems, a variable smoothing algorithm [22] has been extended by [23] to a special version of Problem 1.1 where $\mathfrak{S}$ is a linear operator and $C\coloneqq\mathcal{X}$. The variable smoothing algorithm in [23] uses a smoothed surrogate function $h+{}^{\mu}g\circ\mathfrak{S}$ with the Moreau envelope $^{\mu}g$ of $g$ (see (21)), and its convergence rate, i.e., iteration complexity, $\mathcal{O}(\epsilon^{-3})$ to obtain a certain $\epsilon$-approximate stationary point has been shown under the surjective linearity of $\mathfrak{S}$. This variable smoothing algorithm [23] with its convergence rate analysis has also been extended by [24, 25] to Problem 1.1 where $C$ is a compact embedded submanifold of $\mathcal{X}$, but $\mathfrak{S}$ remains to be linear. In contrast, for variable smoothing algorithms [23, 25] except [24], any asymptotic convergence analysis in terms of a stationary point has not been reported yet, where [24] presents an asymptotic convergence analysis by exploiting the compactness of $C$. Hence, it remains an open issue to establish an asymptotic convergence guarantee of the variable smoothing algorithms for Problem 1.1 even with an unconstrained case $C\coloneqq\mathcal{X}$ or case of a nonlinear differentiable mapping $\mathfrak{S}$.
Contributions of this paper
In this paper, we address the research question: Can we establish a single-loop algorithm for Problem 1.1 with an asymptotic convergence analysis and an iteration complexity of $\mathcal{O}(\epsilon^{-3})$ to obtain an $\epsilon$-approximate stationary point ? To answer affirmatively, we extend a variable smoothing algorithm [23] for Problem 1.1 even with a nonlinear differentiable mapping. Contributions of this paper are three-fold as follows.
  1. (i)
@@
    Parametrization strategy for nonsmooth minimization with manifold constraint. Motivated by parametrization strategies, e.g., [48, 49, 50, 51, 52, 53, 54, 55], we introduce to translate Problem 1.1 into the following unconstrained nonsmooth optimization problem by assuming a manifold constraint set $C$ can be parameterized by a continuously differentiable mapping $\bm{\varphi}:\mathcal{Y}\to C(\subset\mathcal{X})$ defined over a Euclidean space $\mathcal{Y}$ such that $\bm{\varphi}(\mathcal{Y})\coloneqq\{\bm{\varphi}(\bm{y})\in\mathcal{X}\mid\bm{y}\in\mathcal{Y}\}=C$ (see Example 1.3 below for pairs of $C$ and $\bm{\varphi}$).
    Problem 1.2.
    Consider $F=h+g\circ\mathfrak{S}$ and $C\subset\mathcal{X}$ in Problem 1.1. Let $\bm{\varphi}:\mathcal{Y}\to\mathcal{X}$ be a continuously differentiable mapping such that $C=\bm{\varphi}(\mathcal{Y})$, where $\mathcal{Y}$ is a Euclidean space and $\bm{\varphi}$ is called a parametrization of $C$. Then,
@@
      For $\mathcal{X}\coloneqq\mathbb{R}^{2}$, let $C\coloneqq\{\bm{x}\in\mathbb{R}^{2}\mid\left\lVert\bm{x}\right\rVert=1\}$ be the unit circle. Consider $\bm{\varphi}:\mathbb{R}\to\mathbb{R}^{2}:\theta\mapsto[\cos(\theta)\ \sin(\theta)]^{\mathsf{T}}$. Then, $\bm{\varphi}$ is a parametrization of $C$.
@@
      Let $C\subset\mathcal{X}$ be an embedded submanifold of $\mathcal{X}$, endowed with the Riemannian metric induced from $\mathcal{X}$, such that $C$ is connected and closed in $\mathcal{X}$, and let $T_{C}(\bm{x})\subset\mathcal{X}$ be a tangent (vector) space to $C$ at $\bm{x}\in C$ (see (16)). In this case, for any $\bm{x}\in C$, $\mathrm{Exp}_{\bm{x}}:T_{C}(\bm{x})\to C$ with the exponential map [^1] By [32, Defs. 10.6 and 10.7, Thm. 10.8, Exe. 10.12] and [56, Thm. 4.27], the exponential map is uniquely defined over the tangent bundle $TC\coloneqq\{(\bm{x},\bm{v})\mid\bm{x}\in C,\bm{v}\in T_{C}(\bm{x})\}$ as $\mathrm{Exp}:TC\to C:(\bm{x},\bm{v})\mapsto\mathrm{Exp}_{\bm{x}}(\bm{v})\coloneqq\gamma_{\bm{v}}(1)$ with the geodesics $\gamma_{\bm{v}}:I(\subset\mathbb{R})\to C$ (see [32, Def. 5.38][56, (4.16)]) such that $\gamma_{\bm{v}}(0)=\bm{x}$, $\gamma_{\bm{v}}^{\prime}(0)=\bm{v}$ and the geodesic equation $\gamma_{\bm{v}}^{\prime\prime}(t)=0$ is satisfied for all $t\in I$, where $I\subset\mathbb{R}$ is an interval containing $[0,1]$. $\mathrm{Exp}$ is continuously differentiable by [56, Prop. 5.19], and surjective onto $C$ by [56, Cor. 6.21] together with the completeness (closedness in $\mathcal{X}$) of $C$. Hence, $\mathrm{Exp}_{\bm{x}}$ is a parametrization of $C$. See [56, p.136-p.142] for examples of the exponential map of some manifolds.
    In order to validate the idea of parametrization of $C$, we present sufficient conditions, in Theorem 3.2 and Corollary 3.4, for the equivalence between Problem 1.1 and Problem 1.2 in view of finding their stationary points.
@@
    Proposal of a smoothed surrogate function of the nonsmooth function $F\circ\bm{\varphi}$ in Problem 1.2. We introduce a smoothed surrogate function $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ of $F\circ\bm{\varphi}=(h+g\circ\mathfrak{S})\circ\bm{\varphi}$ in Section 4.1 by using the Moreau envelope $^{\mu}g:\mathcal{Z}\to\mathbb{R}\ (\mu\in(0,\eta^{-1}))$ of $g$ (see (21)). In Theorem 4.4 (see also Remark 4.5 (a)), we show that $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ satisfies the gradient consistency property for the original function $F\circ\bm{\varphi}$, where the gradient consistency property will play a key role for establishing an asymptotic convergence property of the proposed algorithm.
    Remark 1.4 (Integral convolution type smoothed surrogate function).
    We can construct another smoothed surrogate function, with gradient consistency property, of $F\circ\bm{\varphi}$ via the integral convolution of $F\circ\bm{\varphi}$ and a kernel function, called mollifier (see [57, Exm. 7.19, Thm. 9.67] or [58, Thm. 4.7]), because $F\circ\bm{\varphi}$ is subdifferentially regular (see Lemma 2.4). However, it is questionable to obtain closed-form expressions of this type of smoothed surrogate functions for $F\circ\bm{\varphi}$ and its gradient, except for simple functions, e.g., the plus function $\mathbb{R}\ni x\mapsto\max\{x,0\}$ [59]. In contrast, the closed-form expressions of our smoothed surrogate function $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ and its gradient are available as computable tools in a case where $g$ is prox-friendly. Moreover, our smoothed surrogate function $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ with a linear $\mathfrak{S}$ and $\bm{\varphi}\coloneqq\mathrm{Id}$ has been employed in the existing variable smoothing algorithms [23, 24, 25], although the gradient consistency property has not been discussed. These situations motivate us to investigate the gradient consistency property of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ in Theorem 4.4.

### v4

@@ Abstract
In this paper, we address a manifold constrained nonsmooth optimization problem involving the composition of a weakly convex function and a smooth mapping under the availability of a parametrization of the manifold. To find a stationary point of the target problem, we propose a variable smoothing-type algorithm by combining the ideas of (i) translating the constrained problem into a Euclidean optimization problem with a parametrization of the constraint set; (ii) exploiting a sequence of smoothed surrogate functions, of the cost function, given with the Moreau envelope of a weakly convex function. The proposed algorithm produces a vector sequence by the gradient descent update of a smoothed surrogate function at each iteration. In a case where the proximity operator of the weakly convex function is available, the proposed algorithm does not require any iterative solver for subproblems therein. By leveraging tools in the variational analysis, we show the so-called gradient consistency property, which is a key ingredient for smoothing-type algorithms, of the smoothed surrogate function used in this paper. Based on the gradient consistency property, we also establish an asymptotic convergence analysis for the proposed algorithm regarding a stationary point. Numerical experiments demonstrate the efficacy of the proposed algorithm.
@@ 1 Introduction
Problem 1.1 has served as key models, e.g., for finding sparse solutions, in wide range of signal processing and machine learning. Such applications include, e.g., sparse Principal Component Analysis (PCA) [11], sparse variable PCA [75], robust sparse PCA [17], orthogonal dictionary learning/robust subspace recovering [24], and sparse spectral clustering [61].
@@
One of simple cases of Problem 1.1 is a convex case where $h$ and $g$ are convex, $\mathfrak{S}$ is linear, and $C\coloneqq\mathcal{X}$. For the convex case, the so-called proximal splitting algorithms have been extensively applied as computationally reliable algorithms (see, e.g., [8, 28]). The algorithms consist of simple steps involving the gradient of $h$ and the so-called proximity operator $\mathrm{prox}_{\mu g}:\mathcal{Z}\to\mathcal{Z}:\widebar{\bm{z}}\mapsto\mathop{\mathrm{argmin}}\limits_{\bm{z}\in\mathcal{Z}}\left(g(\bm{z})+\frac{1}{2\mu}\left\lVert\bm{z}-\widebar{\bm{z}}\right\rVert^{2}\right)$ of $g$ with index $\mu>0$. Although the computation of $\mathrm{prox}_{\mu g}$ requires, in general, solving a convex optimization problem, the proximity operators for many commonly-used functions, e.g., $\ell_{1}$-norm and nuclear norm, have closed-form expressions (see, e.g., [8, 27]).
@@
In Problem 1.1, restricting $g$ to be convex is not always sufficient [23, 71, 85, 1, 86, 52] in order to enhance estimation performance especially for sparsity-aware applications. For example, $\ell_{1}$-norm has been widely used as a best convex approximation of the naive sparsity promoting function, i.e., $\ell_{0}$-pseudonorm, which counts the number of nonzero entries of a vector. However, the use of $\ell_{1}$-norm as $g$ leads to some underestimation of the desired sparse target because $\ell_{1}$-norm is a coercive function (see, e.g., [89]), i.e., $\ell_{1}$-norm excessively penalizes large magnitude of target. To overcome the underestimation effect caused mainly by the sparse promoting convex function, the utilization of weakly convex functions has been attracting great attention [23, 85]. Recent advances in proximal splitting techniques [71, 1, 86, 52] encourage us to find a global minimizer of Problem 1.1 under the convexity of $h+g\circ\mathfrak{S}$ with a linear operator and even with a weakly convex $g$.
Table 1: Summary of existing algorithms for Problem 1.1 and the proposed algorithm, where “cvx” denotes convex function, “w-cvx” denotes weakly convex function, ”lin” denotes linear operator, $\bf C^{1}$” denotes continuously differentiable mapping, $\mathrm{Id}$ denotes the identity operator, and “man” denotes manifold. “Complexity” stands for the first-order oracle complexity for finding an $\epsilon$-approximate stationary point, where $\mathcal{O}$ stands for Landau’s big O-notation.
Assumption Convergence result
Category Method Ref. $g$ $\mathfrak{S}$ $C(\subset\mathcal{X})$ Asymptotic guarantee ? Complexity
Double-loop Proximal gradient [25] cvx lin man# No $\mathcal{O}(\epsilon^{-2})\flat$
Proximal linear [32] cvx $\bf{C^{1}}$ $\mathcal{X}$ No $\mathcal{O}(\epsilon^{-3}\log(\epsilon^{-1}))$
[57] w-cvx $\bf{C^{1}}$ $\mathcal{X}$ Yes -
[80] cvx $\bf{C^{1}}$ man# Yes $\mathcal{O}(\epsilon^{-2})\flat$
[…]
[83] cvx $\bf{C^{1}}$ man No $\mathcal{O}(\epsilon^{-3})$
Single-loop Subgradient [59] w-cvx $\mathrm{Id}$ man# No $\mathcal{O}(\epsilon^{-4})$
Variable smoothing [15] cvx lin $\mathcal{X}$ Yes $\mathcal{O}(\epsilon^{-1})$
[13] w-cvx lin $\mathcal{X}$ No $\mathcal{O}(\epsilon^{-3})$
[10] cvx lin man# Yes $\mathcal{O}(\epsilon^{-3})$
[66] w-cvx lin man# No $\mathcal{O}(\epsilon^{-3})$
Alg. 1 w-cvx $\bf{C^{1}}$ man Yes $\mathcal{O}(\epsilon^{-3})$
@@
    The compactness of $C$ is assumed in [25, 80, 10, 66, 59].
  • $\flat$
    The oracle complexities in [80, 25] are derived in a case where oracle returns the exact solutions of subproblems.
  • $\dagger$}
    The augmented Lagrangian method in [90] can also deal with an additional inequality constraint.
@@
Beyond the convex cases, nonconvex optimization problems with nonconvex constraint have a great potential for flexible formulation to align with real-world tasks. Indeed, the minimization of nonconvex $g\circ\mathfrak{S}$ with a nonlinear smooth mapping $\mathfrak{S}$ is known to have extremely wide applications including, e.g., robust phase retrieval [34], nonnegative matrix factorization [38], robust matrix recovery [20], and robust blind deconvolution [21]. In particular, manifold constrained optimization has been continuingly attracting great attention mainly by its remarkable expressive ability of the target conditions expected to be achieved by a solution (see, e.g., [3, 70, 16, 2]).
@@
Recently, unconstrained nonsmooth optimization algorithms, e.g., [15, 13, 8, 57, 32], have been extended to manifold constrained nonsmooth optimization algorithms [25, 80, 90, 83, 59, 10, 66] (see Table 1). For Problem 1.1, these extended algorithms can be categorized as (i) double-loop algorithms that require an iterative solver to solve a certain subproblem at each iteration (e.g., a proximal gradient method [25], a proximal linear method [80] and augmented Lagrangian methods [90, 83]); (ii) single-loop algorithms without requiring such an iterative solver at each iteration (e.g., a subgradient method [59] and variable smoothing algorithms [10, 66]). Double-loop algorithms [80, 90, 83] can deal with a nonlinear continuously differentiable mapping $\mathfrak{S}$ but with a convex function $g$ in Problem 1.1. An asymptotic convergence guarantee to a stationary point has been shown for [80, 90], while the oracle complexity i.e., the number of oracle calls, for finding an $\epsilon$-approximate stationary point has been shown for [80, 83] (and for [25] under the linear case of $\mathfrak{S}$). However, each update of double-loop algorithms heavily relies on iterative solvers for solving subproblems. Indeed, convergence analyses of double-loop algorithms have been established typically with an exact (or approximate) solution to the subproblem, although such an exact solution is not available in general within a finite number of iterations of an iterative solver. Moreover, an iterative solver for subproblems sometimes takes a lot of time to obtain even an approximate solution to a subproblem (as will be shown in numerical experiments in Section 5.2).
Single-loop algorithms have also been attracting great attention, e.g., [59, 10, 66], for Problem 1.1 because such algorithms do not require solving any subproblem at every iteration. Under the linearity of $\mathfrak{S}$, the oracle complexity for finding an $\epsilon$-approximate stationary point has been shown for [59, 10, 66], where an asymptotic convergence guarantee to a stationary point has been shown for [10] under additional assumptions such as the convexity of $g$ and the compactness of $C$. However, to the best of the authors’ knowledge, for Problem 1.1 with a nonlinear continuously differentiable mapping $\mathfrak{S}$ and a weakly convex function $g$, no single-loop algorithm has been established with a convergence guarantee, in particular an asymptotic convergence guarantee. These situations raise the following research question:
For Problem 1.1 with a manifold constraint $C$, can we design a single-loop algorithm that (i) admits a nonlinear continuously differentiable mapping $\mathfrak{S}$ in the weakly convex composite term $g\circ\mathfrak{S}$, and (ii) converges asymptotically to a stationary point of Problem 1.1 ?
In this paper, we answer this question affirmatively under the availability of a parametrization of $C$ (see Assumption 1.2 below). This assumption is satisfied by a broad class of manifolds encountered in applications (see Example 1.3).
Assumption 1.2 (Parametrization).
@@
The manifold $C$ in Problem 1.1 can be parameterized by a continuously differentiable mapping $\bm{\varphi}:\mathcal{Y}\to C(\subset\mathcal{X})$ defined over a Euclidean space $\mathcal{Y}$, i.e., $\bm{\varphi}(\mathcal{Y})\coloneqq\{\bm{\varphi}(\bm{y})\in\mathcal{X}\mid\bm{y}\in\mathcal{Y}\}=C$. We call $\bm{\varphi}$ a parametrization of $C$ (see Example 1.3 below for pairs of $C$ and $\bm{\varphi}$).
@@
    For $\mathcal{X}\coloneqq\mathbb{R}^{2}$, let $C\coloneqq\{\bm{x}\in\mathbb{R}^{2}\mid\left\lVert\bm{x}\right\rVert=1\}$ be the unit circle. Then, $\bm{\varphi}:\mathbb{R}\to\mathbb{R}^{2}:\theta\mapsto[\cos(\theta)\ \sin(\theta)]^{\mathsf{T}}$ is a parametrization of $C$.
@@
    Consider $C\subset\mathcal{X}$ in Problem 1.1, i.e., $C$ is an embedded submanifold of $\mathcal{X}$, endowed with the Riemannian metric induced from $\mathcal{X}$, such that $C$ is connected and closed in $\mathcal{X}$. Let $T_{C}(\bm{x})\subset\mathcal{X}$ be a tangent (vector) space to $C$ at $\bm{x}\in C$ (see (16)). In this case, for any $\bm{x}\in C$, $\mathrm{Exp}_{\bm{x}}:T_{C}(\bm{x})\to C$ with the exponential map [^1] By [16, Defs. 10.6 and 10.7, Thm. 10.8, Exe. 10.12] and [55, Thm. 4.27], the exponential map is uniquely defined over the tangent bundle $TC\coloneqq\{(\bm{x},\bm{v})\mid\bm{x}\in C,\bm{v}\in T_{C}(\bm{x})\}$ as $\mathrm{Exp}:TC\to C:(\bm{x},\bm{v})\mapsto\mathrm{Exp}_{\bm{x}}(\bm{v})\coloneqq\gamma_{\bm{v}}(1)$ with the geodesics $\gamma_{\bm{v}}:I(\subset\mathbb{R})\to C$ (see [16, Def. 5.38][55, (4.16)]) such that $\gamma_{\bm{v}}(0)=\bm{x}$, $\gamma_{\bm{v}}^{\prime}(0)=\bm{v}$ and the geodesic equation $\gamma_{\bm{v}}^{\prime\prime}(t)=0$ is satisfied for all $t\in I$, where $I\subset\mathbb{R}$ is an interval containing $[0,1]$. $\mathrm{Exp}$ is continuously differentiable by [55, Prop. 5.19], and surjective onto $C$ by [55, Cor. 6.21] together with the completeness (closedness in $\mathcal{X}$) of $C$. Hence, $\mathrm{Exp}_{\bm{x}}$ is a parametrization of $C$. See [55, p.136-p.142] for examples of the exponential map of some manifolds.
With a parametrization $\bm{\varphi}$ of $C$, Problem 1.1 can be translated into the following unconstrained optimization Problem 1.4.
Problem 1.4.
Consider $F=h+g\circ\mathfrak{S}$ and $C\subset\mathcal{X}$ in Problem 1.1 under Assumption 1.2. Then, with a parametrization $\bm{\varphi}$ in Assumption 1.2,
In Section 3, we discuss relations between Problem 1.1 and Problem 1.4. More precisely, we present sufficient conditions, e.g., $N_{C}(\bm{x}^{\star})\supset\mathrm{Ker}((\mathrm{D}\bm{\varphi}(\bm{y}^{\star}))^{*})$ (see (14) for $N_{C}$), or the submersion condition of $\bm{\varphi}$, for the equivalence between Problem 1.1 and Problem 1.4 in view of stationary points (see Theorem 3.2 and Corollary 3.4).
For finding a stationary point of Problem 1.4, we propose a variable smoothing algorithm as an extension of [13] proposed originally for Problem 1.1 in a special case where $C\coloneqq\mathcal{X}$ and $\mathfrak{S}$ is a linear operator (equivalently, Problem 1.4 in a special case where $\bm{\varphi}\coloneqq\mathrm{Id}$ and $\mathfrak{S}$ is a linear operator). We note that the existing variable smoothing algorithm [13] cannot be applied to Problem 1.4 due to the nonlinearity of $\mathfrak{S}\circ\bm{\varphi}$. The proposed algorithm performs a gradient descent update of a smoothed surrogate function $F^{\langle\mu_{n}\rangle}\circ\bm{\varphi}\coloneqq(h+{}^{\mu_{n}}g\circ\mathfrak{S})\circ\bm{\varphi}\ (\mu_{n}\searrow 0)$, of $F\circ\bm{\varphi}$, designed partially with the Moreau envelope $^{\mu}g$ of $g$ (see (21)), where $F^{\langle\mu\rangle}\circ\bm{\varphi}$ is continuously differentiable with $\lim_{\mu\searrow 0}F^{\langle\mu\rangle}\circ\bm{\varphi}(\bm{y})=F\circ\bm{\varphi}(\bm{y})\ (\bm{y}\in\mathcal{Y})$ (see also Lemma 4.1). In a case where $\mathrm{prox}_{\mu g}$ is computable, the proposed algorithm is a single-loop algorithm. To establish an asymptotic convergence analysis of the proposed algorithm, we show the so-called gradient consistency property of $F^{\langle\mu\rangle}\circ\bm{\varphi}$ for $F\circ\bm{\varphi}$ in Theorem 4.4, which makes it possible to express the stationarity condition of Problem 1.4 in terms of a gradient sequence $F^{\langle\mu\rangle}\circ\bm{\varphi}$ with $\mu\searrow 0$. By exploiting the gradient consistency property, we present an asymptotic convergence analysis of the proposed algorithm (see Theorem 4.9). Finally, numerical experiments, in Section 5, demonstrate the efficacy of the proposed algorithm in two scenarios of the sparse spectral clustering and the sparse principal component analysis.
Contributions and Novelties of This Paper
  1. (I)
    A framework of a parametrization strategy for Problem 1.1 via Problem 1.4 is presented in Section 3 where relations between stationarities of Problems 1.1 and Problem 1.4 are presented in Theorem 3.2 and Corollary 3.4. Parametrization strategies have been originally developed for smooth optimization problems (see, e.g., [84, 36, 43, 41, 58, 47, 48, 56]). In contrast, our framework enables us to apply a parametrization strategy even to nonsmooth optimization problems.
@@
    The gradient consistency property of our smoothed surrogate function $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ for $F\circ\bm{\varphi}$ in Problem 1.4 is presented in Section 4.1 (see Theorem 4.4 and Remark 4.5 (a)). The gradient consistency property plays a key role in establishing an asymptotic convergence property of smoothing-type algorithms [26] (see also Appendix A.1). In a simple case where $\bm{\varphi}\coloneqq\mathrm{Id}$ and $\mathfrak{S}$ is a linear operator, our smoothed surrogate function $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ has been used in variable smoothing algorithms [15, 13, 10, 66]; nevertheless, the gradient consistency property of $(h+{}^{\mu}g\circ\mathfrak{S})\circ\bm{\varphi}$ has not been reported yet to the best of the authors’ knowledge. Hence, Theorem 4.4 bridges smoothing-type algorithms, e.g., [26], and variable smoothing algorithms [15, 13, 10, 66].